OpenStudy (anonymous):

Need some help in all H4P2!

OpenStudy (anonymous):

This is how I worked thru it....my given values: VDD=115V, Rs=22orm, RL=120ohm also they give the formula iD=k/2(Vgs-VT)^2 when Vgs>Vt or iD=- when Vgs<Vt 1. so when VI<VT there is a short and VDD=Vout 2.now they give you Vin=6.5 (note Vgs is not equal to Vin b/c of Rs) so to solve ID you solve KVL in the input loop which gives you Vin-IdsRs=Vgs and you plug this into the equation already given...Ids=k/2(Vin-IdsRs-TV)^2 you know K, Rs, Vin and VT so you have to create a quadratic equation to get solve ID you can use http://www.mathwarehouse.com/quadratic/quadratic-formula-calculator.php to solve quadratic ....you will get 2 answers try both *note use many significant figures it will come in handy later 3. To solve Vo use : Vo= Vdd-RL*Id ...we are given Rl and Vdd and we just solved for Id 4. now for delta vo we use VI= Vin + delta vi = 6.5 +1 = 7.5V ....so return to out Ids=k/2(Vin-IdsRs-TV)^2 equation using 7.5V to get our new I ....then plug that into Vo= Vdd-RL*Id ....now subtract this from what you calculated in #3 *note this will be negative and use many significant figures 5.now they give us delta vin=1mV so you can go back and do the work again or you can just take that the bias is linear at this point and the answer should be divided by 1000....but remember they want it in mV so it will be the same as #4 6. just divide what you got in #5 by 1mV and you should get the answer *note #4,#5,#6 are all the same **I used 6 significant digits to get the corrrect answer since when we are using mV the answer needs to be more precise Hope this helps :)

OpenStudy (anonymous):

I was using only 3 decimal places and was getting the wrong answer for #5 and #6....went back and did it with 6 and got it right

OpenStudy (anonymous):

what dd u get for #4

OpenStudy (anonymous):

the answer must be negative

OpenStudy (anonymous):

Thank you so much for the help Cotdamn. Now that's what you call help. Amazing.